OpenStudy (anonymous):

the areas of two similar rectangles are 338 cm^2 and 72 cm^2. what is the similarity ratio of the perimeters?

OpenStudy (dumbcow):

take square root of area ratio 72/338 = 36/169 --> sqrt(36)/sqrt(169) = 6/13